OpenStudy (anonymous):

can u help me with this one One month, a music site observed that 60% of the people who downloaded songs from its site downloaded Q Sam's latest single. The equation below represents this information, where x represents the total number of people who downloaded songs from the site that month. x = 0.6x + 384 How many people who downloaded songs from the site that month downloaded Q Sam's latest single? 576 240 144 960

sam (.sam.):

you need to find x, which is the total number of people who downloaded music that month.

OpenStudy (anonymous):

is it 960 ?

sam (.sam.):

so, x = 0.6x + 384 x - 0.6x = 384 0.4x = 384 0.4x/0.4 = 384/0.4 x = 960 from now you have to look back the question which says that 60 percent of people downloaded this song. then take 60 percent of the value of x. 960 nultiply 0.6 = 576

OpenStudy (anonymous):

can you help me with this one to ? Solve for x: 3(x + 1) = −2(x − 1) + 6 1 4 5 25

sam (.sam.):

3(x + 1) = −2(x − 1) + 6 First multiply out the brackets